Calculate the pH of a 0.0164 M aqueous solution of acetylsalicylic acid (aspirin) (HC9H7O4, Ka = 3.0×10-4).

pH =


Submit Answer
Answer #1

aspirin is weak acid dissociate in water as

HC9H7O4 + H2O   \rightleftharpoons   C9H7O-4 + H3O+

Ka = [C9H7O-4] [H3O+] / [HC9H7O4]

but [C9H7O-4] = [H3O+]

consider [C9H7O-4 ] = [ H3O+] = X

then Ka = [X] [X] / [HC9H7O4]

Ka = X2 / [HC9H7O4 ]

X2 = Ka X [HC9H7O4]

substitute the value

X2 = (3.0 X 10-4) X (0.0164) = 4.92 X 10-6

X = 0.0022181 M

[C9H7O-4 ] = [H3O+] = X = 0.0022181 M

[H3O+] = 0.0022181 M

pH = -log [H3O+] = -log 0.0022181 = 2.65

pH of 0.0164 M HC9H7O4 = 2.65